OpenStudy (anonymous):

Stock prices for each company are given on the table. Prices are given in 8ths of a dollar. The fractions being used are 33 1/2, 67 7/8, and 15 1/4. If someone bought one share of each of these then how much did they spend?

OpenStudy (anonymous):

so u understand mixed fraction

OpenStudy (anonymous):

\[33\frac{1}{2} = 33+ \frac{1}{2} = \frac{66+1}{2} = \frac{67}{2}\]